Hello, the question is: Calculate the average daily balance (in $) for October for a revolving credit account with a previous month's balance of $120 and the following activity. (Round your answer to the nearest cent.)
During October the activities included: October 3rd cash advance of $50.00
October 7th payment of $75.00 October 10th purchase of $29.59 October 16th credit of $30.00 October 25th purchase of 121.60.
Average daily balance =$
I tried to add the previous month's balance and the Purchases and the cash advance while subtracting the credits and the payment. then divided by the number of days in the month (31) and got $6.97 and was wrong.

Found 2 solutions by MathLover1, math_tutor2020:
Answer by MathLover1(20850)
Answer by math_tutor2020(3817)


@MathLover1's solution is incorrect.

Notice she's dividing by 31, but only added up 6 items.
Instead she should be adding up 31 items to represent the 31 days of October

Each day will have a certain balance associated with it. It might help to think of a calendar view as shown in this video here
https://www.youtube.com/watch?v=3TSxhuuLDc0

I do NOT recommend using this method as it's tedious busywork.
There are many repeated values that it would be a pain to write/type out over and over.

-----------------------------------------------

A shortcut is to do a weighted average.

Here's what the chart would look like to show the daily balance for any given period in October.
TimespanNumber of daysBalanceNotes
Oct 1st to Oct 2nd 2120Starting balance
Oct 3rd to Oct 6th4120+50 = 170Cash Advance of $50 (on Oct 3rd)
Oct 7th to Oct 9th3170-75 = 95Payment of $75 (Oct 7th)
Oct 10th to Oct 15th695+29.59 = 124.59Purchase of $29.59 (Oct 10th)
Oct 16th to Oct 24th9124.59-30 = 94.59Credit of $30.00 (happened Oct 16th)
Oct 25th to Oct 31st794.59 + 121.60 = 216.19Purchase of $121.60 (happened Oct 25th)


For the first 2 days of October, the balance is $120 (the previous month's balance)
Then on Oct 3rd, there's a cash advance of $50
This amount is loaned to the customer (it's basically a line of credit)
Therefore, the balance to be repaid increases to 120+50 = 170 dollars. This balance stays in effect from Oct 3rd to Oct 6th as the chart mentions.

On Oct 7th, a payment of $75 is made. This reduces the balance to 170-75 = 95 dollars.
This balance of $95 is in effect from Oct 7th to Oct 9th which is a period of 3 days.

This process is carried out until we arrive at Oct 31st.
Payments and credits reduce the balance; while purchases and cash advances increase the balance.

Here's a simplified version of that same chart. All we really care about is the number of days a certain balance is in effect and the balance for that time period
Number of daysBalance
2120
4170
395
6124.59
994.59
7216.19


A = number of days
B = balance
Form a new column where we multiply A and B
Example: A*B = 2*120 = 240 for the first item in this third column
A = Number of daysB = BalanceA*B
2120240
4170680
395285
6124.59747.54
994.59851.31
7216.191513.33

If you were to add up everything up in the A*B column, then you should get 4317.18

Divide that over the 31 days to get an average daily balance (ADB) of 4317.18/31 = 139.263870967742 = 139.26


Answer: $139.26